# Provably Efficient Online RLHF with One-Pass Reward Modeling

### v2

@@ Appendix F Supporting Lemmas
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have $\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right)$.
@@
[…]
@@ G.1 Implementation Details
  • Online RLHF with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training. Each data point consists of a prompt and two responses with a binary preference label indicating the preferred response. We use the test_prefs split for evaluation.
  • Online RLHF with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies. The global batch size is set to 8 for training. The selection is performed iteratively, where in each iteration, the method selects the most informative samples based on its selection criterion.
@@
0:  Current parameter $\widetilde{\theta}_{t}$, gradient $g_{t}(\widetilde{\theta}_{t})$, learning rate $\eta$, max CG steps $K$, base damping $\lambda_{0}$, error tolerance $\epsilon$
@@
4:     Compute HVP: ~$\widetilde{\mathcal{H}}_{t}p_{k}=\nabla_{\theta}(\nabla_{\theta}\mathcal{L}(\theta)^{\top}p_{k})|_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}p_{k}$
@@
Practical Implementation.   To efficiently implement the OMD update in Eq. (5) without the costly computing and storing the full Hessian matrix, we utilize the Hessian-vector product (HVP) method combined with conjugate gradient descent. The key insight is that the update can be reformulated as solving a linear system. For the OMD update:
$\widetilde{\theta}_{t+1}$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}\approx\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),$
where we ignore the projection operation. The key challenge is computing ~$\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ efficiently. Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$
~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t}),$
where $v$ is the solution we seek. We use the conjugate gradient method with a damped Hessian to solve this system iteratively. The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using automatic differentiation ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda v$, where $\lambda$ is a damping coefficient that ensures numerical stability and $\mathcal{L}(\theta)$ is the loss function. The full algorithm is summarized in Algorithm 5. To ensure stability and convergence, we employ an adaptive damping strategy as $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a damping growth function that can be linear, logarithmic, or cosine. In practice, we set $K=3$ and $\lambda_{0}=0.8$ for all experiments.
@@ G.3 Combined with Adam Optimizer
In previous experiments, we used SGD to update model parameters. In this section, we integrate the methods with the Adam optimizer [Kingma and Ba, 2015], i.e., adding the first and second momentum terms to the model updates. The results, shown in Figure 9, indicate that the Adam optimizer further enhances the performance of our method by leveraging the momentum term to accelerate convergence. With the momentum term, our method remains superior to the MLE-based method; however, the performance gap is reduced. We think that this is because the Adam optimizer incorporates second-order information for optimization, diminishing the advantage of our method compared to the SGD cases.
@@ G.4 Comparison with DPO
We also compare with DPO [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling. To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm. The results are illustrated in Figure 11. While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the methods using our action selection. This result suggests that DPO’s online learning capability remains a challenge. In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that in our deployment stage, a reward model with PPO may be a more suitable choice for sequentially learning from new data.
@@ G.5 Full Update of Reward Model
Figure 11 shows deployment-time adaptation results using the meta-llama/Llama-3.2-1B model, where we update all parameters of the reward model instead of only the final layer. Both our method and MLE use the same action selection strategy. Our approach achieves comparable performance with MLE, indicating that our OMD-based method is still compatible with full-model updates.

### v3

@@ Appendix F Supporting Lemmas
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have
$\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right).$
@@
where $\mathcal{V}\subseteq\mathcal{W}\subseteq\mathbb{R}^{d}$ is a non-empty convex set. Further supposing $\psi(\mathbf{w})$ is 1 -strongly convex w.r.t. a certain norm $\|\cdot\|$ in $\mathcal{W}$, then there exists a $\mathbf{g}_{t}^{\prime}\in\partial\ell_{t}\left(\mathbf{w}_{t+1}\right)$ such that
@@ G.1 Implementation Details
  • Passive data collection: We randomly choose $30,000$ samples from the UltraFeedback-binarized dataset’s train_prefs split for training. Each sample consists of a prompt and two responses with a label indicating the preferred response. We use the test_prefs split for evaluation.
  • Active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies. The global batch size is set to 8 for training. The selection is performed iteratively, where in each iteration, the method selects the most informative samples based on its selection criterion.
@@
0: Current parameter $\widetilde{\theta}_{t}$, gradient $g_{t}(\widetilde{\theta}_{t})$, learning rate $\eta$, max CG steps $K$, parameter $\lambda_{0}$, $\epsilon$
@@
4:  Compute HVP: $\widetilde{\mathcal{H}}_{t}p_{k}=\nabla_{\theta}(\nabla_{\theta}\mathcal{L}(\theta)^{\top}p_{k})|_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}p_{k}$
@@
Update details.   As described in Section 6.1, we can implement the OMD update using the HVP with conjugate gradient descent. The full algorithm is summarized in Algorithm 5. In our experiments, we set $K=3$ and $\lambda_{0}=0.8$ and choose the linear function $f(t/T)=t/T$ as the damping function.
@@ G.3 Combined with Adam Optimizer
In previous experiments, we used SGD to update model parameters. In this section, we integrate the methods with the Adam optimizer [Kingma and Ba, 2015], i.e., adding the first and second momentum terms to the model updates. The results, shown in Figure 5, indicate that the Adam optimizer further enhances the performance of our method by leveraging the momentum term to accelerate convergence. With the momentum term, our method remains superior to the MLE-based method; however, the performance gap is reduced. This may be because the Adam optimizer incorporates second-order information for optimization, diminishing the advantage of our method compared to the SGD cases.
@@ G.4 Comparison with DPO
We also compare with DPO [Rafailov et al., 2023] in the deployment stage. As a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling. To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm. The results are illustrated in Figure 7. While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the methods using our action selection. This result suggests that DPO’s online learning capability remains a challenge. In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that a reward model with PPO may be a more suitable choice for sequentially learning from new data.
@@ G.5 Full Update of Reward Model
Figure 7 shows deployment-time adaptation results using the Llama-3.2-1B model, where we update all parameters of the reward model instead of only the final layer. Both our method and MLE use the same action selection strategy. Our approach achieves comparable performance with MLE, indicating that our OMD-based update method is still compatible with full-model updates.
G.6 More Foundation Models and Datasets
In this section, we provide more experimental results about other foundation models and datasets.
Figure 8 shows the training and evaluation curves for reward model learning under passive data collection using the Qwen2.5-7B-Instruct model. We compare our method with MLE and report the loss and accuracy over training. Our method consistently shows stable training dynamics and competitive evaluation performance compared to MLE, suggesting its effectiveness in offline settings.
Figure 10 present results for online RLHF with active data collection using the same Qwen model. Figure 9(a) shows training loss curves, while Figure 9(b) reports evaluation accuracy over training iterations. Table 9(c) further compares various methods (Rand-MLE, Active-MLE, Rand-OMD, and our approach) in terms of final accuracy and training time. While Active-MLE achieves slightly higher accuracy, our method provides significant speedup in training time with comparable performance, highlighting the efficiency of our approach.
Figure 10 illustrates the deployment-time performance of various methods on the Ultrafeedback dataset. We split the dataset into 20 chunks and measure cumulative rewards across these chunks. Our method demonstrates robust adaptation capabilities, achieving competitive reward accumulation.
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 8: For Qwen2.5-7B-Instruct model with passive data collection, we compare our method with MLE. We report average accuracy and loss curve of the reward model.
(a) training loss
(b) evaluation accuracy
Method ACC (%) Time (s)
Rand-MLE 67.36 $\pm$ 0.5 4653 $\pm$ 121
Active-MLE 67.25 $\pm$ 0.4 4701 $\pm$ 103
Rand-OMD 66.80 $\pm$ 0.5 1312 $\pm$ 47 0
Ours 67.11 $\pm$ 0.5 1325 $\pm$ 54 0
(c) training time
Figure 9: For Qwen2.5-7B-Instruct with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) evaluation accuracy and training time.
Figure 10: Results of deployment-time adaptation for Qwen2.5-7B-Instruct.
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 11: For online RLHF with passive data collection on the Llama-3-8B-Instruct model on the Mixture2 dataset, we compare our method with MLE. We report average accuracy and loss curve of the reward model.
Finally, Figure 11 shows results on the Llama-3-8B-Instruct model trained on the Mixture2 dataset in a passive data collection setup. Similar to earlier observations, our method achieves competitive or superior performance compared to MLE, both in terms of training and evaluation loss/accuracy, demonstrating its generality across different model and dataset combinations.
Appendix H Broader Impact
Our work advances the efficiency of RLHF, a central technique in aligning large language models with human values and preferences. By proposing a new one-pass reward modeling method that eliminates the need to store historical data and re-train from scratch, we reduce the computational and environmental costs commonly associated with online RLHF pipelines. This could enable the development and deployment of aligned language models by institutions with limited resources.
However, the broader deployment of RLHF, particularly in an online and adaptive setting, raises important ethical and societal considerations. On the positive side, it can enable more responsive and value-aligned AI systems, with potential applications in education, healthcare, and accessibility. Yet, the ability to iteratively adapt to user feedback in deployment may also increase the risk of reinforcing harmful biases or being gamed by adversarial users, especially in high-stakes or open-ended domains.